OpenStudy (anonymous):

Reduction formula.. ∫sin^5x cos^2x dx

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

sin^5x = sinx * sin^4x = sinx*(1-cos^2x)^2 helps ?

OpenStudy (anonymous):

then what do i have to use substitution formula?

OpenStudy (anonymous):

first expand you will have 3 terms (3 integrals) in each of them you will use the substitution (u=cosx)

OpenStudy (anonymous):

note that in each of them you will have cosx to some power multiplied by sinx since -sin is the derivative of cosx the integrals are easy

OpenStudy (anonymous):

okay i am gonna do it and then will write the solution to u??u are here for how many hours?

OpenStudy (anonymous):

go for it.. even if i wont be here im sure someone else will be able to check your work

OpenStudy (anonymous):

okay thank u.. :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

∫sin^5x cos^2x dx let cosx=z =>d/dx cosx=dz/dx =>-sinx dx=dz =>sinx dx=-dz Now, ∫sin^5x cos^2x dx=∫sinx.sin^4x cos^2x dx =∫sinx(1-cos^2x)^2. cos^2x dx =∫sinx(1-2cos^2x+cos^4x) cos^2x dx =∫sinx.cos^2x dx-∫sinx 2cos^2x cos^2x dx+∫sinx cos^4x cos^2x dx =∫-z^2 dz-∫-2z^4 dz+∫-z^8 dz =-z^2+2/2+2+2 z^4+4/4+4-z^8+8/8+8+c =-z^4/4+2z^8/8-z^16/16+c =-cos^4x/4+2cos^8/8-cos^16/16+c @Coolsector

OpenStudy (anonymous):

oh god it will be cos^12/12 sorry

OpenStudy (anonymous):

@eliassaab :( plz help

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

\[I=\int\limits \sin ^{5}x \cos ^{2}x~dx=\int\limits \sin ^{4}x \sin x \cos ^{2}x~dx\] \[=\int\limits \left( 1-\cos ^{2}x \right)^{2}\cos ^{2}x \sin x~dx=\int\limits \left( 1-2\cos ^{2} x+\cos ^{4}x \right)\cos ^{2}x \sin x~dx\] \[=\int\limits \left( \cos ^{2}x-2\cos ^{4}x+\cos ^{6}x \right)dx\] \[=\frac{ -\cos ^{3}x }{ 3 }+\frac{2 \cos ^{5}x }{ 5 }-\frac{ \cos ^{7}x }{ 7 }\]

OpenStudy (anonymous):

correction write sin x near dx outside the bracket in last integral line.

OpenStudy (anonymous):

oh sorry i forgot to use the formula its x^n+1/n+1

OpenStudy (anonymous):

My head is quite hot..it's tempting !!